OpenStudy (anonymous):

simplify 12x^5 over 20x^2

OpenStudy (anonymous):

12x^3/20 3/5x^3

OpenStudy (anonymous):

\[\frac{ 12x^5 }{ 20x^2 }=\frac{ 3*4 x^{5-2}}{ 5*4 }=\frac{ 3x^3 }{ 5 }\]

OpenStudy (anonymous):

thank you again (:

OpenStudy (anonymous):

np babe ;)